OpenStudy (anonymous):

What is the length of the longest side of a triangle that has the vertices (-3, 2), (3, 2), and (3, 6)? A. 5 units B. 2√ ̅13 units C. 2√ ̅15 units D. √ ̅5 units

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

 If you can see the diagram, it shows the approximation of the three points. The lines between them form a right triangle, since two of the points are on y=2 and two points are on x=3. So, in a right triangle the longest side has to be the hypotenuse, which is between (3, 6) and(-3,2). To calculate the distance between the two, you can either use the distance formula, or since this is a right triangle, the Pythagorean theorem. The Pythagorean theorem is easier, it goes like this: a^2 + b^2 = c^2 a and b are the two shorter legs and c is the hypotenuse. so (from the diagram) : 4^2+6^2= c^2 16+36=c^2 52=c^2 \[\sqrt{52}\]= \[\sqrt{13*4}\]= \[2\sqrt{13}\] That's it! Hope that helps!